### v2

# PRISM-Consult: A Panel-of-Experts Architecture for Clinician-Aligned Diagnosis

Abstract
We present PRISM-Consult, a clinician-aligned panel-of-experts architecture that extends the compact PRISM sequence model into a routed family of domain specialists. Episodes are tokenized as structured clinical events; a light-weight router reads the first few tokens and dispatches to specialist models (Cardiac–Vascular, Pulmonary, Gastro–Oesophageal, Musculoskeletal, Psychogenic). Each specialist inherits PRISM’s small transformer backbone and token template, enabling parameter efficiency and interpretability. This initial study evaluates a scoped panel of five specialist families defined by high-impact ED diagnostic groups. On real-world Emergency Department cohorts, specialists exhibit smooth convergence with low development perplexities across domains, while the router achieves high routing quality and large compute savings versus consult-all under a safety-first policy. We detail the data methodology (initial vs. conclusive ICD-9 families), routing thresholds and calibration, and report per-domain results to avoid dominance by common events. The framework provides a practical path to safe, auditable, and low-latency consult at scale, and we outline validation steps—external/temporal replication, asymmetric life-threat thresholds, and multi-label arbitration—to meet prospective clinical deployment standards.
I Background
Clinical care unfolds as time-ordered sequences of events that include presenting complaints, assessing vitals, diagnostic orders, laboratory observations, and evolving diagnoses. The core challenge is to turn the earliest, sparsest portion of this sequence into reliable predictions that (1) anticipate likely next events and differentials and (2) support time-critical decision-making. Our goal is to achieve this with models that are fast, auditable, and robust to noisy early signals. In the Emergency Department (ED), where minutes matter and presentations are often ambiguous, we represent encounters as tokenized event streams and learn to predict the next event and, by extension, the patient’s diagnostic trajectory.
We previously introduced (PRISM) a compact, domain-focused approach to sequence modeling[1]. We focused on the ED setting as a proving ground for sequence-based clinical decision support, emphasizing early-prefix prediction that can surface high-value differentials and guide downstream diagnostic actions. PRISM represented Emergency Department (ED) episodes as tokenized event streams and trained a small decoder-only transformer to predict the next event and, by extension, the evolving diagnostic trajectory [1]. PRISM emphasized low-latency, interpretable modeling by (i) restricting the event schema to diagnostics, labs, and diagnoses (excluding procedures and medications), (ii) employing a hand-designed token template (e.g., [DIAG]_ICD9_410.xx, [OBS]_LAB_TROP:HIGH, [ACTION]_ORD_ECG), (iii) capping sequence length (e.g., 512 tokens) and vocabulary size (compact custom lexicon), and (iv) utilizing a moderate-capacity backbone (e.g., 6 layers, $d_{model}{=}256$) suitable for clinical deployment constraints.
Initially applied to chestpain presentations in the ED, PRISM demonstrated strong calibration and efficient inference, validating the premise that carefully scoped tokenization and model size can yield clinically useful predictions without the cost or opacity of very large general-purpose models. However, the same design choices that improve efficiency, namely a streamlined vocabulary and a single-target clinical domain, also narrow the applicability of the model: real-world emergency department presentations often traverse multiple organ systems, and differential diagnoses for a single symptom (e.g., chest pain) span cardiac, pulmonary, gastro-oesophageal, musculoskeletal, and psychogenic causes. This motivates an extension from a single-domain backbone to a routed panel of domain specialists.
II INTRODUCTION
II-A Clinical motivation and use case
In the Emergency Department (ED), clinicians must make time-critical decisions from sparse, noisy early signals: a chief complaint, a small number of initial vitals, and the first diagnostic actions (e.g., ECG, labs, imaging orders). At this stage, the task is rarely “predict the final diagnosis” in a single step; instead, clinicians form an evolving differential and decide which actions are most urgent and informative. This creates a practical need for early decision support that can suggest likely next steps and plausible differentials while remaining fast, auditable, and conservative under uncertainty.
II-B Why clinician-aligned routing
Real ED care is organized around consult services and organ-system expertise (e.g., cardiology, pulmonary/critical care, gastroenterology). We therefore frame early prediction as a routing problem: from the earliest events, decide which specialist perspectives are most appropriate to engage. This design supports safety-first behavior (e.g., consult multiple specialists when uncertain) and enables explicit control over compute/latency trade-offs while preserving clinical interpretability: a routed specialist corresponds to a clinically recognizable “who should weigh in?” decision.
II-C Why transformer specialists rather than a single monolith
A single sequence model can be effective, but ED trajectories often cross domains and contain heterogeneous patterns (e.g., dyspnea spanning pulmonary embolism, pneumonia, heart failure, anxiety). Specialization allows each model to learn domain-consistent thresholds, priors, and event semantics while retaining a shared tokenization template and backbone for parameter efficiency. The resulting panel improves coverage without requiring an ever-expanding single-model vocabulary, and it allows domain-specific calibration and evaluation aligned to clinical consequences.
In this paper we present PRISM-Consult, a routed panel-of-experts architecture that extends the PRISM methodology from a single-domain model to a family of specialist models orchestrated by a lightweight router. The router interprets the earliest events (symptoms and first diagnostic cues) and dispatches the episode to one or more specialist models aligned with clinical organ systems. Each specialist inherits the PRISM tokenization template, shares embeddings for consistency, and is fine-tuned on domain-filtered cohorts labeled by definitive ICD-9 families. This preserves PRISM’s efficiency and interpretability while expanding clinical coverage.
Our contributions are threefold. First, we formalize a clinical routing surface based on initial ICD-9 symptom codes and map it to specialist targets defined by conclusive diagnostic families. Second, we implement a minimalist router and parameter-efficient specialist adapters that retain PRISM’s compact footprint while providing cross-domain accuracy. Third, we report partial results across cardiac, pulmonary, gastro-oesophageal, musculoskeletal, and psychogenic domains, showing smooth convergence and low development perplexities (near $\sim$2.0 for all five completed specialists), thereby validating that the PRISM backbone generalizes when routed to domain-consistent corpora.
II-D Why a Panel of Experts? Tokenization Trade-offs and Clinical Coverage
Its tokenization strategy sits at the core of PRISM’s efficiency. A more specific token inventory (e.g., separate tokens for finely binned labs, highly granular diagnoses) captures nuanced patterns and can improve in-domain accuracy and calibration. Yet specificity increases vocabulary size, fragmenting data across rare tokens and slowing inference. Conversely, a more generic inventory (broader bins, fewer code variants) compacts the vocabulary and speeds inference but risks losing discriminative power across heterogeneous diseases.
This creates a fundamental tension in single-model designs: optimizing the token schema for one domain (e.g., cardiology) may degrade performance in another (e.g., pulmonary embolism vs. pneumothorax), where different signals and thresholds matter.
II-E Heterogeneity, Comorbidity, and Cross-Specialty Trajectories in the ED
Further complicating diagnostic predictions is the fact that emergency presentations are intrinsically heterogeneous and frequently comorbid: a single complaint (e.g., chest pain, dyspnea, syncope) often admits plausible explanations across multiple organ systems, and early measurements (vitals, first labs, first orders) carry overlapping, sometimes contradictory signals. In MIMIC-IV ED cohorts, the same prefix of events can legitimately evolve toward cardiac (AMI, dissection), pulmonary (PE, pneumothorax, pneumonia), gastro–oesophageal (Boerhaave, spasm, reflux), musculoskeletal, or psychogenic end points, or some combination thereof, and patients routinely carry chronic conditions (CKD, COPD, diabetes, atrial fibrillation) that shift priors and confound intermediate tests. These realities create diagnostic pathways that traverse traditional service boundaries and unfold via cross-specialty handoffs.
This cross-disciplinary drift introduces several modeling hazards for single-domain specialists. First, boundary undercoverage: a specialist tuned to one domain’s vocabulary and thresholds may systematically down-weight rare but life-threatening out-of-domain hypotheses that share early features (e.g., PE vs. AMI in troponin-positive dyspnea). Second, calibration shift: domain-specific likelihoods calibrated in-domain can become miscalibrated when the latent etiology changes mid-episode, especially under sparse, prefix-only supervision. Third, premature closure: strong local patterns (e.g., ECG orders $\rightarrow$ ischemia) can anchor the trajectory before counter-signals appear (e.g., pleuritic descriptors, D-dimer, CTA), yielding myopic next-event proposals.
Sequential structure further compounds these risks. Clinical events arrive as partially ordered sets (co-occurring orders, batched lab panels) with irregular time gaps, and many ED timelines are better viewed as sequences of sets than pure token strings. Methods that enforce set-level invariance or augment code streams with ontology signal (e.g., CCS/knowledge-graph embeddings) help preserve meaning across co-occurrences and synonyms, but they do not by themselves resolve cross-service drift when the generator of observations changes during workup[32] [33]. Likewise, zero-shot generative rollouts over patient-health timelines demonstrate breadth but still reflect the chosen pretraining mix and may blur service-specific decision thresholds [29].
From a probabilistic standpoint, the ED setting resembles large-context inference over many interacting latent causes. Classic graphical approaches (e.g., QMR-DT) formalize comorbidity and conditional dependence but become computationally brittle at realistic scales [34]. Autoregressive transformers approximate these high-order conditionals efficiently, yet as already noted above, a single backbone must compromise between compact vocabularies (speed, stability) and domain-specific expressivity (rare tokens, narrow thresholds). In practice, optimizing tokenization and priors for one service can degrade discrimination in another, particularly near safety-critical boundaries.
These observations motivate a routed, multi-disciplinary engagement approach. Concretely it: (i) reads the earliest prefix and estimate calibrated probabilities over clinical domains; (ii) consults one or more domain specialists when warranted (with asymmetric, safety-first thresholds for life-threatening etiologies); (iii) arbitrates multi-label suggestions with deterministic, auditable policies; and (iv) fails open when confidence is low or vitals breach guardrails. Such conditional computation aligns with mixture-of-experts and expert-routing advances—activating only the parameters relevant to the emerging hypothesis space while preserving latency and traceability.
PRISM-Consult is designed with this approach in mind. By holding the token template constant—preserving interpretability and shared embeddings—while routing episodes to domain-specialized fine-tunings. Each specialist selectively expands or emphasizes a local subset of tokens relevant to its domain (e.g., pleural processes, gas exchange markers for pulmonary; ischemic markers and ECG patterns for cardiac), without forcing the global vocabulary to balloon. The router thus acts as a clinical analogue of referral: “right expert, right time,” mitigating the specificity–coverage trade-off inherent to a single global model.
III Related Work
III-A Diagnostic Progressions and LLM Token Prediction
In large language models (LLMs), each token—representing a word or sub-word unit—has clear pairwise probabilities relative to other tokens. For example, given a token such as ”New,” the subsequent token might have high probability for ”York,” ”Zealand,” or ”Jersey.” Individually, these pairwise probabilities are straightforward. However, as the sequence of tokens expands—considering an entire preceding sentence or paragraph—the combinatorial complexity and contextual dependencies make it impossible to precisely calculate probabilities using first order correlations alone. Hence, deep learning and attention mechanisms in transformer architectures (like GPT) approximate these complex conditional dependencies, capturing long-range contexts and intricate relationships between tokens.
By analogy, diagnostic medicine faces a similar complexity. Individual diagnostic tests may exhibit well-defined pairwise probabilities for specific diseases—just as pairwise token probabilities can be well-established. However, when attempting to predict the next best test or disease diagnosis given the entire preceding sequence of events (the patient’s complete history, including medical tests, symptoms, prior diagnoses, lifestyle factors, and comorbidities), the scenario becomes profoundly more complex. The interactions between these elements are intricate, context-dependent, and frequently nonlinear, making exact probabilistic inference practically infeasible.
Thus, Bayesian statistical approaches quickly reach limitations. Transformer-based diagnostic models, analogous to language models, can approximate the complex contextual dependencies inherent in patient histories, making them potentially more capable of accurately predicting subsequent medical events (tests, results, or diagnoses). These models implicitly learn from data to reflect patient-level complexity more naturally, offering both predictive capability and interpretable attention scores to highlight influential factors in diagnostic reasoning.
III-B Prior Work in Transformer and Deep Learning Models for Sequential Clinical Diagnosis
In contrast to high-level summations of pre-existing efforts, which are considerable and highly varied in nature, we opt to focus in on prior work of immediate relevance to our proposed approach, by zeroing in on a methodology-specific approaches that have informed our own experimental design, and serve as a basis for which our work attempts to build on.
III-B1 Tokenization and Representation of Medical Data
The first step in adapting transformer models to diagnostic-sequence prediction is to convert a patient’s richly structured history into a sequence of discrete tokens. Early work relied on standardised vocabularies such as ICD-9/ICD-10, CPT and Read codes, treating each code as an individual token [30, 31]. A coarser alternative aggregates thousands of raw codes into a few hundred clinically coherent groups, reducing sequence length while retaining clinical meaning [30].
More granular pipelines tokenise free-text concepts, symptoms and history fragments extracted from clinical narratives [28]. The ETHOS framework extend this idea further by encompassing: admissions, diagnoses (ICD-10-CM), procedures (ICD-10-PCS), medications (ATC), laboratory results (LOINC) and even inter-event time gaps. Under the ETHOS framework, these events are mapped to 1–7 tokens each, then ordered chronologically into a patient-health timeline (PHT) [29].
Because multiple events can occur simultaneously during a visit, several authors represent a patient record as a sequence of sets, forcing the model to be order-invariant within each set. DPSS is the canonical example of this strategy [32]. Finally, semantic enrichment with external knowledge graphs (or hierarchical ontologies such as CCS) can be injected by concatenating concept embeddings with ontology embeddings, as done in the SETOR framework [33].
III-B2 Transformer Model Architectures in Diagnostic Prediction
Traditionally model choice varied depending on the prediction task. Encoder-only families (BERT, ALBERT, MedAlbert) excel at classification problems such as early disease detection from longitudinal code streams; MedAlbert, for instance, uses a 6-layer ALBERT encoder to flag incipient lung cancer three years in advance [30]. Decoder-only designs (GPT-like) suit generative forecasting: ETHOS, for instance, employs a causal decoder to roll out future PHTs in a zero-shot setting [29]. Hybrid encoder–decoder stacks (e.g., TransformEHR) translate past encounters into predicted future ICD sequences [31].
Domain-aware modifications are common. SETOR feeds visit-level embeddings— augmented with CCS-derived ontology vectors and continuous-time positional encodings—into a multi-layer transformer encoder dubbed the Patient Journey Transformer [33]. Such customisations can improve both performance and interpretability without abandoning the core attention machinery of the original architecture.
III-C Additional Clinical foundation models and domain adaptation
Foundation models trained on biomedical and clinical corpora underpin many recent advances in clinical NLP. Early domain-adapted encoders such as BioBERT [2] and ClinicalBERT [3] demonstrated that continued pretraining on PubMed/PMC and EHR notes yields substantial gains for NER, RE, QA, and risk prediction. PubMedBERT [4] further showed that training from scratch on biomedical text can outperform continual-pretraining approaches on multiple benchmarks. Generative clinical LMs have since emerged: BioGPT [5] brought GPT-style pretraining to biomedical literature, while GatorTron and GatorTronGPT scaled to billions of parameters using mixed corpora that include large EHR collections, improving a wide array of downstream tasks [6, 7]. Complementing these efforts, the Med-PaLM line explored instruction tuning and medical QA evaluation at scale, reporting expert-competitive answers on MultiMedQA and related evaluations [8, 9].
Despite strong aggregate performance, deploying a single generalist model for all problems can be suboptimal in clinical operations. Differences in vocabulary, labeling practices, and decision latencies across services (e.g., cardiology vs. psychiatry) motivate domain-aligned specializations with small, auditable backbones, provided we can direct cases reliably to the right specialist. This motivates the routed panel-of-experts framing adopted in PRISM-Consult.
III-D Mixture-of-experts and routing among specialists
Mixture-of-Experts (MoE) and conditional computation architectures enable input-dependent parameter activation and have been central to efficient scaling. GShard [10] and Switch Transformer [11] demonstrated that sparse expert routing can train trillion-parameter class models with near-constant per-token cost, provided stability and load-balancing constraints are met. Subsequent work proposed alternative routing rules (e.g., Expert Choice) to improve load balancing and efficiency [12]. In parallel, “routing among models” at inference—choosing which pre-trained system to run for a given query—has become an active area, with routers trained from preference data or heuristics to trade off cost, quality, and latency [13, 14].
Beyond general LLM systems, ensemble/agent-style methods explicitly coordinate multiple specialists. Recent “mixture-of-agents” approaches show that structured collaboration among models can improve reasoning and robustness [15]. In healthcare, most prior work focuses on single-model evaluation (e.g., LLMs for triage or diagnosis) [16, 17] or traditional clinical decision support (CDS) [18]. There remains a gap for lightweight, auditable routers that triage early clinical prefixes to compact, domain-specific experts, optimizing safety and compute while preserving traceability—precisely the niche that PRISM-Consult targets.
III-E Clinical sequence similarity and pathway clustering
A long line of work in clinical sequential pattern mining and pathway analysis focuses on defining similarity between patient trajectories and clustering them into clinically meaningful pathways. Temporal similarity measures are central in this setting because clinical workflows depend not only on event identity but also on ordering, repetition, and time gaps [36]. Pathway clustering methods adapt classical alignment/edit-distance algorithms (e.g., Needleman–Wunsch variants) to group patient journeys into interpretable clusters [37]. Related studies also examine clustering diagnosis-code sequences jointly with demographics to discover stratified pathway structure [39]. These approaches primarily target cohort discovery, summarization, and retrospective pathway characterization; by contrast, PRISM-Consult targets prospective early-prefix prediction and clinician-aligned dispatch to specialist models. Nonetheless, similarity-based pathway work informs what constitutes a “near miss” versus a “severe miss,” motivating our cost-sensitive evaluation.
III-F Probabilistic sequence models and stochastic automata
Classical probabilistic approaches such as Markov models, hidden Markov models, and stochastic automata have been applied to encoded clinical sequences due to their interpretability and tractability [41]. These methods can perform well when the state space is limited and the temporal process is relatively stationary. However, ED event streams are heterogeneous and high-dimensional, with long-range dependencies and sparse early prefixes that are difficult to capture with low-order Markov assumptions without aggressive state aggregation. We therefore focus on compact transformer specialists with shared tokenization, while explicitly comparing against and discussing probabilistic alternatives where appropriate.
III-G Positioning PRISM-Consult
PRISM-Consult operationalizes a routed specialist architecture tailored to ED presentations: (i) compact specialists (same PRISM backbone and schema) trained on domain-consistent corpora; (ii) a calibrated router that reads the first $K$ tokens and dispatches to one or more experts under safety-first thresholds; and (iii) per-domain evaluation and macro-averaging to avoid dominance by common presentations. Compared with monolithic clinical LMs, this design aligns with operational realities (service-specific vocabulary and SLAs), supports explainability (specialist audit trails and calibrated probabilities), and can scale horizontally by adding experts or refining routing policies, drawing on MoE and LLM-routing principles while remaining deployable in resource-limited clinical settings.
IV Methods
IV-A Overview
PRISM-Consult extends the original PRISM framework—a compact, domain-focused sequence model over structured clinical events—into a routed, multi-specialist system. All data handling, event schema design, tokenization, and baseline modeling choices follow PRISM; below we restate those elements in full and then describe the extensions for the router and specialist models. Throughout, we reference the prior work for provenance [1] while ensuring this section is self-contained.
IV-B Data and Preprocessing
IV-B1 Data Source
Data for this study was extracted from the MIMIC-IV database, an extensive electronic health records repository collected from patients admitted to the Beth Israel Deaconess Medical Center (BIDMC) between 2008 and 2019. The database encompasses detailed clinical information, including demographics, vital signs, laboratory test results, diagnostic procedures, and discharge diagnoses captured during hospital stays [19].
IV-B2 Data structure and cohort window
Each record for this model is comprised of a comprehensive longitudinal record of a patient’s diagnostic journey, across clinical episodes. Episodes begin at admission and end at patient discharge.
Timelines consist of timestamped structured events drawn from:
  1. 1.
    Patient Information (E.g., age, gender)
  2. 2.
    Admission and Discharge Events
  3. 3.
    Diagnostics and orders (e.g., ECG ordered, CTA chest ordered).
  4. 4.
    Laboratory observations (e.g., troponin value, D-dimer positive/negative).
  5. 5.
    Diagnoses encoded in ICD-9-CM.
Procedures and medication administrations are intentionally excluded, mirroring PRISM’s emphasis on a concise and clinically interpretable vocabulary focused on diagnostic reasoning. All events are normalized to a patient-relative clock (Initial admission =0=0) and sorted chronologically and then in a structured manner detailed below to ensure consistency in batched event tokenization.
IV-B3 Inclusion/exclusion Criteria
Our study included all adult ED encounters with at least one qualifying initial (pre-diagnosis) symptom code recorded during the index window, along with at least one of the identified terminal diagnoses, in order to have a comprehensive set of patients traversing the diagnostic progression across clinical disciplines.
IV-B4 Initial (Pre-Diagnosis) ICD-9-CM Code Set — Router Inputs
Figure 1 details the set of preliminary diagnostic codes a patient presents with that are notionally indicative of a diagnosis pathway. Inclusion of at least one such code is a requisite for inclusion in the study.
Figure 1: Initial set of diagnoses, notionally tied to associated diagnostic grouping
IV-B5 Final (Conclusive) ICD-9-CM Code Set — Specialist Gold Labels
Figure 2 details the set of final, or target, Gold-Label diagnostic codes representing an ultimate diagnostic determination by a clinical expert. Patients must ultimately be diagnosed with one of these conditions for inclusion in the study cohort.
Figure 2: Set of final ’Gold Label’ diagnostic codes, representing an ultimate diagnostic determination
Cohort composition note: The cohort is defined by the study’s previously delineated gold-label diagnostic families rather than the unconditional ED case mix; as a result, domain prevalence reflects selection criteria and may over-represent certain high-acuity presentations (e.g., cardiopulmonary). To mitigate dominance by any single domain, we report per-domain results in addition to macro-averaged metrics, and we treat evaluation on broader, less-selected cohorts as an external validation objective.
IV-B6 Sequence construction
Once eligible patients were determined, their medical histories were extracted and converted to token sequences using a fixed event-precedence policy, based foremost on chronological timing and then a predetermined hierarchy of token types: DIAG >> LAB >> ORDER, with explicit time-gap markers inserted between events when inter-event intervals exceed pre-specified thresholds (e.g., 1 hour, 6 hours) to preserve chronology. A final alphabetical ordering within each token-type is then done to ensure that batched sequences are ordered consistently across episodes (this ensures consistency in prediction accuracy given the singular ’next token’ generation schema to the PRISM model). Each sequence is truncated or windowed to a maximum length of 512 tokens for an entire patient chronology, potentially spanning multiple admission episodes. The final (gold-label) diagnosis token is not included on the input side for training tasks that predict it, preventing label leakage.
IV-B7 Tokenization template and vocabulary
We use a compact, hand-designed token schema with type prefixes to preserve semantics: [^1] These templates are inherited from PRISM and reused verbatim. New domains add tokens but do not alter templates.
[…]
  • Special sentinels: [BOS], [EOS], [PAD], [UNK].
Vocabulary growth is controlled by (i) whitelisting code families relevant to target presentations and (ii) merging infrequent variants into [UNK] or higher-level bins (e.g., rare lab subtypes). This yields a compact vocabulary that supports low-latency inference while maintaining clinical interpretability.
IV-C Model backbone (PRISM) Configuration and Training
The baseline PRISM model is a decoder-only transformer with:
  • L=6L{=}6 transformer blocks; model dimension $d_{model}{=}256$; MLP expansion $4d$; $n_{heads}{=}4$.
  • Learned absolute positional embeddings; tied input/output token embeddings for parameter efficiency.
  • Dropout $0.1$ in attention and MLP layers; layer normalization pre-attention.
The objective is next-token prediction over the event sequence (autoregressive cross-entropy). For temporal awareness, an auxiliary time-to-next-event head (Huber loss) can be added during pre-training; this head is discarded at inference.
IV-C1 Optimization
We train with AdamW (weight decay $0.01$), linear warmup over the first 5% of steps followed by cosine decay to 10% of the peak learning rate. Typical settings: batch size 64–128 sequences, peak LR in $1\text{e-}3$ to $2\text{e-}4$ depending on corpus size, gradient clipping at 1.0, early stopping on dev NDCG@3. Mixed-precision training is used when supported.
IV-C2 Evaluation (PRISM baseline)
Primary metrics for model performance were Top-1/Top-3 next-event recall, NDCG@k, and calibration (Brier score, reliability curves). For diagnosis-prediction ablations, AUROC/PR for specific gold-label families were measured. All metrics are computed per domain and macro-averaged to avoid dominance by common events.
IV-D PRISM-Consult (panel of experts) Configuration and Training
Early ED prefix (first $K$ events) Tokenization + optional time features Lightweight calibrated router $p(d\mid\text{prefix})$ over domains Dispatch policy Top-1 / Top-2 specialists Fail-open under low confidence Safety guardrails: low confidence or danger thresholds $\Rightarrow$ consult-all Domain specialists PRISM backbone + LoRA adapters Domain-aligned next-event / diagnosis-family scores Deterministic arbitration / merge (e.g., Cardiac >> Pulmonary >> …) Outputs Predictions + uncertainty Audit log: routed domains, scores, thresholds
Figure 3: PRISM-Consult end-to-end routing pipeline.
PRISM-Consult’s framework extends PRISM by the expansion of the following elements:
  1. 1.
    A router trained on the earliest events of each episode to emit one or more domain flags (Cardiac–Vascular, Pulmonary, Gastro–Oesophageal, Musculoskeletal, Psychogenic). Inputs are the first 2–5 coded events represented as bag-of-concepts TF–IDF features projected to a 256-dim space, optionally concatenated with simple temporal features (e.g., time since triage). The router is a logistic-regression classifier and a Platt scaling router with sigmoid outputs and class weighting to up-weight life-threatening domains.
  2. 2.
    A set of specialist models—one per domain—initialized from the PRISM backbone and fine-tuned on domain-filtered corpora defined by the final gold-label codes specified above. To preserve parameter sharing and inference speed, each specialist uses low-rank adaptation (LoRA) on attention and feed-forward layers while keeping shared embeddings frozen; adapter ranks and learning rates are selected on a held-out dev set per domain.
  3. 3.
    A dispatch policy at inference: if any life-threatening domain (e.g., AMI, PE, dissection) exceeds a calibrated threshold, a single high-priority specialist is invoked; otherwise the top-2 domains are consulted in parallel and their suggestions are merged by a deterministic arbitration layer (Cardiac >> Pulmonary >> Gastro >> Musculoskeletal >> Psychogenic).
IV-D1 Labeling surface (initial & final codes)
For each episode, only codes occurring before the first definitive diagnosis are considered “initial” to prevent leakage; when multiple definitive diagnoses are present (e.g., AMI + PE), multi-label targets are assigned.
IV-D2 Calibration, safety, and auditability
All classifier outputs (router and specialists) were temperature-calibrated on a development fold using cross-entropy minimization. We logged router logits, specialist logits, and the final arbitration decision per episode to enable post hoc review. If the router emits uniformly low confidence (all domain probabilities below a safety threshold) or vitals cross pre-defined danger thresholds, the system fails open to parallel consultation of all specialists.
IV-D3 Compute and reproducibility
Experiments for the router were run on modern GPUs using deterministic seeds. Data preprocessing and training pipelines were version-controlled; model checkpoints, tokenizer vocabularies, and ICD-9 codebooks (both initial and final sets) are archived with SHA checksums.
Reproducibility and artifacts. All experiments were run on NVIDIA A100 (Ampere) GPU with 40GB VRAM, with with an 8-core CPU and 83.5 GB of system RAM) and software stack Linux Ubuntu, CUDA 12.x, NVIDIA proprietary driver.
We provide the full training and evaluation pipeline, including preprocessing scripts, tokenizer vocabulary construction, fixed random seeds, and configuration files. An accompanying public repository includes a minimal runnable notebook that reproduces key preprocessing and metric computations on a small illustrative example, enabling validation of the reported methodology.
Code availability. Code for PRISM-Consult, including preprocessing, training, and evaluation scripts, along with a minimal runnable example notebook, is available in the PRISM repository. [^3] https://github.com/lmlevine/prism
IV-D4 Overview of Light-Weight Router Design
We model early Emergency Department (ED) episodes as a prefix time-series classification task. After the first $K$ coded events (default K=5K{=}5), the router produces calibrated probabilities over five specialist domains: Cardiac–Vascular, Pulmonary, Gastro–Oesophageal, Musculoskeletal, and Psychogenic. Episodes may be multi-label in principle, though the present corpus yielded disjoint (single-label) episodes. The router dispatches to one or more PRISM-Consult specialists under a safety-first policy.
IV-D5 Data ingestion and harmonization
Tokenized episodes reside in five domain directories (Cardiac, Pulmonary, Gastro–Oesophageal, Musculoskeletal, Psychogenic).
Each record is mapped to a canonical schema:
$\text{episode}\;=\;\bigl(\texttt{episode\_id},\;\mathbf{e}=[e_{1},\ldots,e_{L}],\;\texttt{time\_feats}\in\mathbb{R}^{\leq 2}\bigr),$
where optional time_feats contain simple scalars (minutes-to-first-order; max inter-event gap). Episodes with the same episode_id across directories are merged (longest token list retained; non-empty time_feats preferred), and a 5-dim multi-hot label $\mathbf{y}\in\{0,1\}^{5}$ indicates domain membership. In this study, directories were disjoint by construction, so $\|\mathbf{y}\|_{0}{=}1$.
IV-D6 Proportional sampling
Let $\mathcal{E}=\{(\mathbf{e}_{i},\mathbf{y}_{i})\}_{i=1}^{N}$ denote the merged episode table. To obtain a target size $T$ while preserving domain mixture, we compute per-domain prevalence $p_{d}$ on $\mathcal{E}$ and allocate quotas $q_{d}\approx p_{d}T$. We first fill quotas with domain-exclusive episodes, then top up from remaining sets (including potential multi-label cases), avoiding duplicates; if fewer than $T$ remain after deduplication, we top up uniformly at random. Setting T=0T{=}0 disables sampling.
IV-D7 Prefix expansion (anytime supervision)
To enable predictions after each early event, we expand episodes into prefixes of lengths $\ell\in\{1,\dots,\min(K,L_{i})\}$:
$\mathcal{P}\;=\;\bigl\{(i,\ell,\;\mathbf{e}_{i,1:\ell},\;\mathbf{y}_{i},\;w_{\ell}=\ell/K)\bigr\}.$
Each prefix inherits the final label $\mathbf{y}_{i}$ but only uses the first $\ell$ tokens as input, preventing label leakage. We optionally use $w_{\ell}$ as a sample weight to mildly emphasize later (more informative) prefixes.
IV-D8 Featurization: TF–IDF $\rightarrow$ SVD
We join prefix tokens with spaces to form a short “document” and compute 1–2-gram TF–IDF with min_df=2. A 256-dim truncated SVD yields a compact vector $\mathbf{z}_{i,\ell}\in\mathbb{R}^{256}$; optional time_feats ($\leq$2 scalars) are concatenated to form $\tilde{\mathbf{z}}_{i,\ell}$.
IV-D9 Router model and calibration
[…]
IV-D10 Routing policy and threshold tuning
At inference, with $\hat{\mathbf{p}}=\{\hat{p}_{d}\}_{d=1}^{5}$:
  1. 1.
    If a life-threatening domain (Cardiac, Pulmonary) satisfies $\hat{p}_{d}\geq\tau_{\text{hi}}$, route top-1 (the argmax).
  2. 2.
    Else if $\max_{d}\hat{p}_{d}\geq\tau_{\text{lo}}$, route top-2.
  3. 3.
    Else, fail-open to all five experts.
[…]
IV-D11 Safety behavior
If calibrated probabilities are uniformly low (i.e., $\max_{d}\hat{p}_{d}{<}0.25$) or if triage vitals cross pre-defined danger thresholds, the router fails open to all experts and emits an audit record (raw logits, thresholds, selected route). Temperature scaling is re-checked on each model refresh to maintain probability fidelity.
IV-D12 Evaluation
Discrimination. ROC-AUC and PR-AUC are reported per domain on the held-out test set using calibrated probabilities.
Routing quality
With routed set $R_{i}$ and truth $Y_{i}$ for episode $i$:
$\text{Recall}_{\text{any}}=\tfrac{1}{|\mathcal{I}|}\sum_{i\in\mathcal{I}}\mathbf{1}[R_{i}\cap Y_{i}\neq\varnothing],\quad$
$\text{Recall}_{\text{all}}=\tfrac{1}{|\mathcal{I}|}\sum_{i\in\mathcal{I}}\mathbf{1}[Y_{i}\subseteq R_{i}].$
Life-threat recall is computed on:
$\{i:Y_{i}\cap\{\text{Cardiac},\text{Pulmonary}\}\neq\varnothing\}$
and requires $R_{i}$ to include at least one of these domains.
Compute proxy and latency
We report $\mathbb{E}[|R|]=\tfrac{1}{N}\sum_{i}|R_{i}|$ and an estimated latency $L_{i}=L_{\text{router}}+\sum_{d\in R_{i}}L_{d}$ using fixed per-expert times.
Anytime curves
Metrics are stratified by prefix length $\ell=1,\dots,K$ to quantify earliness.
Anytime performance
To quantify earliness, we repeat the above at each prefix length $\ell\in\{1,\dots,K\}$, plotting discrimination and routing recalls as functions of $\ell$.
IV-D13 Ablations and robustness checks
We assess: (i) $K\in\{2,3,5\}$; (ii) text-only vs. text+time features; (iii) uncalibrated vs. calibrated probabilities; and (iv) alternative linear heads (linear SVM with Platt). As sanity baselines we compare against consult-all, fixed Cardiac+Pulmonary, and a single generalist PRISM variant (no routing).
V Results
V-A Specialist Model Training & Performance
V-A1 Cohorts and training setup
A total of N=20,436N{=}20{,}436 Emergency Department (ED) episodes met all inclusion criteria for PRISM-Consult. Domain-eligible cohorts are not disjoint (multi-label episodes may appear in multiple domain pools). Where noted, some domains were scoped to a capped training subset to ensure compute parity and class balance. All specialists inherit the PRISM backbone and tokenization; optimization and early-stopping follow the protocol in the Methods section. Unless otherwise stated, training used the default causal language modeling loss of autoregressive cross-entropy) with mixed precision.
V-A2 Domain-specific outcomes (Specialist Model Set Performance)
Each specialist exhibited stable convergence with steadily improving validation loss across epochs, consistent with the baseline PRISM behavior on cardiac timelines. Table 4 reports the best development loss, derived perplexity ($\mathrm{PPL}=\exp(\mathrm{loss})$), and percentage reduction from epoch 1 to the best epoch.
Figure 4: Development-set performance by specialist model. Perplexity is $\exp(\text{Val Loss})$. $\Delta$ Val Loss is the percentage reduction from epoch 1 to the best epoch. Scoped domains used capped training sets for balance.
Cardiac–Vascular (AMI/UA/pericardial/aortic/HCM)
Validation loss decreased from $1.8397$ at epoch 1 to $0.7917$ at epoch 5 ($56.97\%$; $\mathrm{PPL}{=}2.21$), closely mirroring the original PRISM cardiac behavior and supporting the transferability of the backbone.
Pulmonary (PE/pneumothorax/pleura/pneumonia)
Validation loss improved from $1.4753$ to $0.7041$ by epoch 5 ($52.27\%$; $\mathrm{PPL}{=}2.02$), indicating a well-captured structure in early pulmonary presentations.
Gastro–Oesophageal (GERD/spasm/Boerhaave/ulcer perforation)
On a scoped training pool of 4,500 episodes, validation loss fell from $1.1303$ to $0.7004$ ($38.03\%$; $\mathrm{PPL}{=}2.01$) by epoch 5, with steady epoch-on-epoch gains and no signs of overfitting.
Musculoskeletal (costochondritis/rib injury/chest-wall contusion)
Despite the smallest cohort (N=523N{=}523), validation loss dropped from $2.6927$ to $1.2692$ ($52.87\%$; $\mathrm{PPL}{=}3.56$). The higher perplexity is consistent with data scarcity and etiologic heterogeneity; nevertheless, the relative improvement and stable trajectory indicate learnability under limited data.
Psychogenic (panic/anxiety/hyperventilation)
Training on a scoped subset of 4,500 episodes converged from $0.8910$ to $0.6289$ ($29.42\%$; $\mathrm{PPL}{=}1.88$), the lowest perplexity among the specialists, reflecting a relatively compact symptom-to-diagnosis mapping for this domain.
Figure 5: Cross-specialist Training Loss over epochs for both training and validation cohorts
V-A3 Cross-domain interpretation
Figure 4 visualizes the training of all specialist models across epochs. All five specialists converge to low development perplexities (three near ${\sim}2.0$ and one below $2.0$), with smooth validation curves and substantial loss reductions from epoch 1 (Table 4). This cross-disciplinary consistency—achieved without altering tokenization, model size, or optimization hyperparameters—validates our central design choice: the compact PRISM backbone, when routed to domain-consistent corpora, yields strong and stable learning dynamics beyond its original cardiac scope. The musculoskeletal model’s higher perplexity likely reflects limited sample size and broader label variance; we anticipate improvements with targeted data augmentation and modest adapter-rank increases.
V-A4 Training curves (dev) and calibration
For each specialist, validation loss decreased monotonically through epoch 5 with no divergence from training loss, suggesting adequate regularization (dropout 0.1; early stopping at epoch 5). Final checkpoints correspond to the minima reported above. Temperature scaling for probabilistic outputs is fit on the development fold, with reliability plots to be included in the appendix.
V-B Router Training Results
V-B1 Cohort and training rows
Across domain directories we ingested N=13,801N{=}13{,}801 unique episodes: Cardiac 1,1281{,}128 (8.2%), Pulmonary 3,1503{,}150 (22.8%), Gastro–Oesophageal 4,5004{,}500 (32.6%), Musculoskeletal $523$ (3.8%), Psychogenic 4,5004{,}500 (32.6%). Episodes were single-label by design (no cross-directory duplicates). Prefix expansion with K=5K{=}5 yielded 69,00569{,}005 training/evaluation rows.
V-C Cohort and Routing Policy Summary
Figure 6: Light-Weight Router Training Cohort and Overall Results Summary
V-C1 Selected thresholds and routing mix
Fig 5. details specific thresholds and overall results mix. The development grid selected $(\tau_{\text{hi}},\tau_{\text{lo}})=(0.70,\,0.30)$ under our objective. With these thresholds, the router predominantly returned top-1 or top-2 routes; fail-open was rare. The expected consulted experts per episode on test was $1.565$, a $\sim\!69\%$ reduction versus consult-all (5 experts).
V-D Discrimination (per-domain; macro-averaged)
Fig. 6 details domain specific descriminitive results, along with macro-averaged results across domains for both the development and testing cohorts.
Figure 7: Domain-specific discriminitive results
V-D1 Safety and routing quality
[…]
Figure 8: End-to-end PRISM-Consult system summary (routing safety and efficiency)
VI Discussion
VI-A Cross-disciplinary performance of PRISM
PRISM-Consult extends the original PRISM design from a cardiology-focused study to a routed, multi-specialist system while preserving the compact backbone and token template. Empirically, all five specialists trained with the same optimization recipe and achieved smooth, monotonic declines in development loss with low final perplexities (near $\sim$2.0 for Cardiac, Pulmonary, Gastro–Oesophageal, and Psychogenic; higher but improving for Musculoskeletal). This cross-disciplinary consistency, obtained without increasing model width/depth or altering tokenization, indicates that the PRISM backbone transfers effectively when exposed to domain-consistent corpora and calibrated with light routing. That is, a single representational space and schema can support heterogeneous organ-system reasoning as long as data partitioning and supervision align with clinical provenance.
Additionally the router’s near-perfect Recallany and high Recallall further suggest that early tokens carry enough discriminative signal to select the appropriate specialist(s) reliably, even under strict latency constraints.
VI-B Cost-sensitive clinical utility (future work)
While we report standard discrimination and ranking metrics, real clinical deployment is cost-sensitive: the consequence of a miss depends on severity (e.g., failing to identify a time-critical cardiopulmonary presentation) and the consequence of “over-consulting” is often primarily resource or latency cost. A principled extension is to report an expected-utility (or expected-cost) metric defined over the routed specialist set. For episode $i$ with true domain set $Y_{i}$ and routed set $R_{i}$, one can define:
$C_{i}=\sum_{d\in Y_{i}\setminus R_{i}}\lambda_{d}\;+\;\gamma\sum_{d\in R_{i}\setminus Y_{i}}1,$ (1)
where $\lambda_{d}$ is higher for life-threatening domains and $\gamma$ captures the comparatively lower cost of over-routing. This yields an explicit compute–safety Pareto curve as router thresholds vary. In the current study, we did not retain per-episode routed-set outputs for all runs, which prevents retrospective computation of such utility-based metrics from aggregate results alone. We therefore include this evaluation as a priority direction for future work and as a recommended standard for high-stakes clinical sequence prediction.
VI-C Next steps: validation and extension
We outline three directions to harden and extend this framework:
  1. 1.
    External and temporal validation. Replicate training/evaluation on (i) a second site with distinct coding habits and lab panels; (ii) a temporally held-out slice to assess drift.
  2. 2.
    Safety-first routing refinements. Enforce a hard development constraint of life-threatening recall $\geq 0.98$ via (a) asymmetric thresholds (lower $\tau_{\mathrm{hi}}$ for Cardiac/Pulmonary than for other domains), (b) a guardrail that includes both Cardiac and Pulmonary whenever $\max(\hat{p}_{\mathrm{card}},\hat{p}_{\mathrm{pulm}})\geq\tau_{\mathrm{lo}}^{\mathrm{life}}$, and (c) per-head isotonic calibration when under-confidence is detected near decision boundaries. Re-report the compute–safety Pareto frontier (expected experts vs. life-threat recall).
  3. 3.
    Broader coverage and richer inputs. Add specialists (e.g., aortic catastrophes vs. general vascular, neurologic mimics) as new gold-label families mature; incorporate small, domain-agnostic time features (minutes-to-first-order; max inter-event gap) that are already supported by the router. Where multi-label episodes exist (e.g., AMI+PE), explicitly train/evaluate multi-label routing quality (Recallall at top-$k$) and measure arbitration behavior.
Operationally, we recommend prefix-stratified audits (metrics by $\ell=1..K$), route-mix telemetry (top-1/top-2/fail-open), and episode-level audit logs (router/specialist logits, thresholds, decision) to support prospective QA and IRB-facing safety reviews.
VI-D Limits
Three significant limitations temper interpretation:
Multi-label handling (limitation and evaluation plan)
First, the present corpus is effectively single-label across directories, so the multi-label routing regime was not strongly exercised. We therefore treat multi-label arbitration as a priority validation objective. In future work we will evaluate confirmed multi-diagnosis cases (e.g., AMI+PE) and report multi-label routing quality (e.g., Recallall@k / top-$k$ coverage, calibration by prefix length), along with structured audits of arbitration outcomes and fail-open behavior in safety-critical presentations.
Limited Clinical Domains and Skewed Patient-Mix
Second, this study evaluates only five specialist families selected to support clear gold-labeling and clinically meaningful ED coverage; this does not represent the full breadth of ED consult services. In addition, cohort composition reflects the selected diagnostic families and may be skewed toward higher-acuity presentations resulting in class imbalance (e.g., the relatively small Musculoskeletal pool) likely contributes to higher perplexity and calibration variance for that head; targeted sampling, modest adapter rank increases, and data augmentation (token normalization for near-synonymous events) are straightforward mitigations for class imbalance, and future work should also aim to expand the specialist panel (e.g., neurologic, infectious/sepsis, renal/metabolic, toxicologic/trauma) and evaluate scalability of routing and arbitration as the panel grows, and (ii) validate generalization on broader, less-selected cohorts and across sites/time periods to better reflect real-world ED case mix.
Fixed Positional Embeddings may Prove too Rigid for Model Adaptability
Third, learned absolute positional embeddings fix the maximum context length and may not extrapolate; if longer horizons are needed, rotary or ALiBi encodings can be substituted with minimal disruption to the rest of the stack.
Additionally, we note that threshold selection in the current run favored a near-constraint point (life-threat recall <0.98<0.98 on test); future reports should hard-enforce prespecified safety constraints and present confidence intervals via bootstrap over episodes.
Finally, to better contextualize model accuracy and efficiency, PRISM-Consult should be benchmarked against (i) classical classifiers (e.g., XGBoost) trained on the same early-prefix feature representation, and (ii) compact encoder models (e.g., ClinicalBERT/BioBERT-style) trained to predict domain/diagnosis from prefix tokens.
VII Conclusion
PRISM-Consult operationalizes a clinician-aligned panel-of-experts by pairing a calibrated, light-weight router with parameter-efficient PRISM specialists trained on domain-consistent corpora. Using only the earliest tokens, the router attains high coverage and substantial compute savings relative to consult-all, while specialists exhibit smooth, low-perplexity convergence across disparate organ systems—evidence that the PRISM backbone generalizes beyond its original cardiac scope. With minor policy and calibration refinements to meet strict life-threat recall targets, the framework provides an auditable, low-latency pathway to deployable clinical decision support that routes each episode to the right expert at the right time.